# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

### v1

@@ E.2 Computational Costs
Table 4: Comparison of training time per epoch, epochs to convergence, and sampling time (for 1,000 forecast series over a 30-step horizon) across Tree, Diamond, and Layer structures.
@@
E.3 Synthetic Data Experiments
@@
DoFlow $\mathbf{0.61_{\pm.13}}$ $\mathbf{0.62_{\pm.13}}$ $\mathbf{0.69_{\pm.15}}$ $\mathbf{0.69_{\pm.16}}$ $\mathbf{0.74_{\pm.17}}$ $0.75_{\pm.17}$
@@
E.4 Hydropower System
@@
In Figure 11, $X_{1}$ and $X_{2}$ denote the horizontal and vertical rotational vibrations of the hydraulic turbine, which directly drive the generator’s operation. The generator’s horizontal and vertical dynamics are represented by $X_{3}$ and $X_{4}$, while $X_{5}$ captures the generator’s current output delivered to the transformer ($X_{8}$) for voltage regulation and transmission to the power grid. The Metering and Control (M&C) unit ($X_{7}$) continuously monitors the generator’s performance and turbine stability to ensure coordinated operation. In addition, $X_{6}$ represents an auxiliary generator state (thermal signal) recorded independently for diagnostic purposes but not directly connected in the causal structure.
Table 7 reports the RMSE results for observational and interventional time-series forecasting in the hydropower system. For interventional forecasting, we use 12 real power outages where the root nodes Turbine $X_{1}$ and $X_{2}$ fail, causing the entire system to an outage. In this setting, the root nodes are treated as intervened by the breakdown signals. The reported averages and standard deviations are computed over 12 runs with batch size $B=1$ for the interventional case, in contrast to 50 runs with batch size $B=128$ for the observational case.
@@
Figure 10 presents two segments of DoFlow log-probabilities during outage time windows. In each case, minute 50 marks the actual outage event. Notably, DoFlow’s log-probability output becomes abnormal well before the outage occurs—sometimes as early as 20 minutes prior (e.g., $X_{6,t}$ in the right panel) and as late as 10 minutes prior. This allows anomalies to be detected in advance of the true outage.
E.5 Cancer Treatment Effects
@@
During the forecasting window (days 56566262), the treatment schedule is replaced by one of ten pre-defined intervention plans $\mathcal{I}^{j}=\{(i,t):X_{i,t}\leftarrow m_{i,t}^{j}\}$, where $j$ indicates the $j$-th treatment plan. At each time step $t$, DoFlow estimates the interventional tumor outcome by sampling from the learned flow model:
@@
Table 8 reports the normalized RMSE results for causal treatment effect estimation on cancer tumor outcomes. We compare our method with three established baselines: CRN (Counterfactual Recurrent Network) [5], RMSN (Recurrent Marginal Structural Network) [35], and MSM (Marginal Structural Model) [43]. The datasets follow the construction in Bica et al. [5], where $\gamma_{c}$ and $\gamma_{r}$ denote the treatment-application budgets for chemotherapy and radiotherapy, respectively. Some baseline results in Table 8 are adopted directly from Bica et al. [5].

### v2

@@ F.1 Computational Costs
Table 3: Comparison of training time per epoch, epochs to convergence, and observational sampling time (for 1,000 forecast series over a 30-step horizon) across Tree, Diamond, and Layer structures.
@@
F.2 Synthetic Data Experiments
@@
Table 5: CRPS for observational and interventional time–series forecasting across causal structures: Tree, Diamond, FC-Layer, and Chain. Reported values are averaged over 50 test batches, each containing 128 test series.
Tree Diamond FC-Layer Chain
Additive NLNA Additive NLNA Additive NLNA Additive NLNA
DoFlow $0.22_{\pm.01}$ $\mathbf{0.21_{\pm.01}}$ $\mathbf{0.27_{\pm.02}}$ $\mathbf{0.29_{\pm.02}}$ $\mathbf{0.13_{\pm.01}}$ $\mathbf{0.15_{\pm.01}}$ $\mathbf{0.17_{\pm.01}}$ $\mathbf{0.20_{\pm.02}}$ $\mathbf{0.11_{\pm.01}}$ $\mathbf{0.14_{\pm.01}}$ $\mathbf{0.16_{\pm.02}}$ $\mathbf{0.14_{\pm.02}}$ $\mathbf{0.24_{\pm.02}}$ $\mathbf{0.25_{\pm.02}}$ $\mathbf{0.38_{\pm.04}}$ $\mathbf{0.39_{\pm.04}}$
DeepVAR $0.24_{\pm.02}$ $0.27_{\pm.02}$ $0.32_{\pm.03}$ $0.35_{\pm.03}$ $0.16_{\pm.01}$ $0.24_{\pm.02}$ $0.21_{\pm.02}$ $0.26_{\pm.03}$ $0.14_{\pm.01}$ $0.19_{\pm.02}$ $0.20_{\pm.02}$ $0.25_{\pm.03}$ $0.27_{\pm.02}$ $0.29_{\pm.02}$ $0.44_{\pm.05}$ $0.47_{\pm.06}$
MQF2 $\mathbf{0.19_{\pm.02}}$ $0.29_{\pm.02}$ $0.34_{\pm.03}$ $0.40_{\pm.04}$ $0.17_{\pm.02}$ $0.26_{\pm.03}$ $0.23_{\pm.02}$ $0.31_{\pm.04}$ $0.13_{\pm.01}$ $1_{\pm.02} 0.22±.0$$2_{\pm.02}$ $0.29_{\pm.03}$ $0.28_{\pm.03}$ $0.32_{\pm.03}$ $0.46_{\pm.05}$ $0.52_{\pm.07}$
@@
DoFlow $\mathbf{0.61_{\pm.13}}$ $\mathbf{0.66_{\pm.13}}$ $\mathbf{0.30_{\pm.06}}$ $\mathbf{0.69_{\pm.16}}$ $\mathbf{0.71_{\pm.17}}$ $\mathbf{0.41_{\pm.07}}$
@@
Figure 9: Chain” (NLNA) counterfactual forecasting results. Node $X_{1,t}$ is intervened. DoFlow provides a single forecast in green; the orange lines indicate the true counterfactual future.
F.3 Hydropower System
@@
Reliable modeling and uncertainty quantification are increasingly central to modern signal processing [Dong et al., 2020, Nemani et al., 2023, Gao et al., 2023, Wu et al., 2024a]. Here, we highlight a new perspective enabled by DoFlow, causal reasoning in time series, through a hydropower signal processing case study.
In Figure 10, $X_{1}$ and $X_{2}$ denote the horizontal and vertical rotational vibrations of the hydraulic turbine, which directly drive the generator’s operation. $X_{3}$ represents an auxiliary turbine state that is affected by $X_{1}$ and $X_{2}$. The generator’s horizontal and vertical dynamics are represented by $X_{4}$ and $X_{5}$, while $X_{6}$ captures the generator’s current output delivered to the transformer ($X_{8}$) for voltage regulation and transmission to the power grid. The Metering and Control (M&C) unit ($X_{7}$) continuously monitors the turbine’s performance and stability to ensure coordinated operation.
Table 7 reports the RMSE results for observational and interventional time-series forecasting in the hydropower system. For interventional forecasting, we use 12 real power outages where the root nodes $X_{1}$ and $X_{2}$ fail, causing the entire system to an outage. In this setting, the root nodes are treated as intervened by the breakdown signals. The reported averages and standard deviations are computed over 12 runs with batch size $B=1$ for the interventional case, in contrast to 50 runs with batch size $B=128$ for the observational case.
@@
For anomaly detection, we use DoFlow as a likelihood-based scoring model over predicted future windows. Fix a node $i$ and a current time $t$ in the test sequence. We treat the observed past $\{\mathbf{X}_{1},\dots,\mathbf{X}_{t}\}$ as context, sample a future window of length $L=30$ minutes from DoFlow, $\{\hat{X}_{i,t+1},\dots,\hat{X}_{i,t+L}\}$, and compute the conditional log-density that DoFlow assigns to this generated forecast trajectory (i.e., the model-implied confidence of its own prediction) using the change-of-variables formula. Formally, this yields the node-wise anomaly score
$\ell_{i}(t):=\log p_{\theta,X_{i,t+1:t+L}}\!\left(\hat{x}_{i,t+1:t+L}\mid\hat{H}_{i,t}\right),$
where $\hat{x}_{i,t+1:t+L}$ denotes the forecast window decoded from sampled base noise $z_{i,t+1:t+L}$ under DoFlow conditioned on $\hat{H}_{i,t}$. Here, the log-likelihood is computed using the Eq. (13), restricted to the window [t+1,t+L][t{+}1,t{+}L]:
$\log p_{\theta,X_{i,t+1:t+L}}\!\left(\hat{x}_{i,t+1:t+L}\mid\hat{H}_{i,t}\right)=\sum_{u=t+1}^{t+L}\left[\log q\!\left(z_{i,u}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{i,\theta}\bigl(x_{i,u}(s),s;\,\hat{H}_{i,u-1}\bigr)\,ds\right],$ (59)
with $\hat{x}_{i,u}=\Phi_{\theta}^{-1}(z_{i,u};\hat{H}_{i,u-1})$ and $\hat{H}_{i,u-1}$ the RNN summary of the past up to u1u{-}1 as defined in Section 2.3. Intuitively, $\ell_{i}(t)$ measures how typical the entire upcoming 30-minute trajectory of node $i$ appears under the dynamics learned from normal-operation data by the model.
To obtain a quantitative decision rule, we first compute $\ell_{i}(t)$ on all test windows that do not overlap any outages and collect the empirical distribution of these scores for each node $i$. Let $\eta_{i}$ denote the 10th percentile of this normal-score distribution. We then flag time $t$ as locally abnormal for node $i$ whenever
$\ell_{i}(t)<\eta_{i}.$
To avoid reacting to isolated noise spikes, we declare an anomaly only if at least one node exhibits $5$ consecutive minutes of low scores, i.e., there exists $i$ and a run of times $t,t{+}1,\dots,t{+}4$ such that $\ell_{i}(\tau)<\eta_{i}$ for all $\tau$ in that run. The anomaly detection time for a given outage is the earliest $t$ at which this criterion is met; in Figure 11, the corresponding detection markers are plotted relative to the known outage time (minute 50 in each panel).
In Figure 11, notably, DoFlow’s log-probability output becomes abnormal well before the outage occurs—sometimes as early as 20 minutes prior (e.g., $X_{6,t}$ in the right panel) and as late as 10 minutes prior. This allows anomalies to be detected in advance of the true outage.
F.4 Cancer Treatment Effects
@@
Notably, unlike our other experiments, here only the tumor-volume node $Y_{t}$ is modeled with a CNF. The treatment actions $\mathbf{X}_{t}$ (chemotherapy and radiotherapy assignments/dosages) are fully observed and directly provided at every time step. We still use RNNs to summarize the joint history of past treatments $\mathbf{X}_{t}$ and past tumor outcomes $Y_{t}$ into a hidden state $\hat{H}_{t-1}$, which serves as the conditioning variable for the CNF of $Y_{t}$.
During the forecasting window (days 56566262), the treatment schedule is replaced by one of ten pre-defined intervention plans $\mathcal{I}^{j}=\{(i,t):X_{i,t}\leftarrow m_{i,t}^{j}\}$, where $j$ indicates the $j$-th treatment plan and $j\in\{1,2,\dots,10\}$. At each time step $t$, DoFlow estimates the interventional tumor outcome by sampling from the learned flow model:
@@
Table 8 reports the normalized RMSE results for causal treatment effect estimation on cancer tumor outcomes. We compare our method with three established baselines: CRN (Counterfactual Recurrent Network) Bica et al. [2020a], RMSN (Recurrent Marginal Structural Network) Lim [2018], and MSM (Marginal Structural Model) Mansournia et al. [2012]. The datasets follow the construction in Bica et al. [2020a], where $\gamma_{c}$ and $\gamma_{r}$ denote the treatment-application budgets for chemotherapy and radiotherapy, respectively. Some baseline results in Table 8 are adopted directly from Bica et al. [2020a].